OpenStudy (anonymous):

Find the exact value of sin82.5

OpenStudy (kira_yamato):

Express the above as 90 - x first

OpenStudy (kira_yamato):

Apply the half-angle formula twice

OpenStudy (dumbcow):

\[\sin(90-x) = \cos x\] \[\cos (\frac{x}{2}) = \sqrt{\frac{\cos x +1}{2}}\]

OpenStudy (anonymous):

is it 7.5 degrees?

OpenStudy (anonymous):

90-82.5?

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

0.73018385915316689224111517182195006273281159789180934210

OpenStudy (anonymous):

to be exact....

OpenStudy (dumbcow):

then 30/2 = 15 and 15/2 = 7.5

OpenStudy (dumbcow):

@RICARDOismyfirstname , that is a decimal approximation :) you cant express irrational number exactly with deciamals

OpenStudy (anonymous):

cos\[\cos(7.5\div2)=\sqrt{cosx+1}\]

OpenStudy (anonymous):

Yeah, man you're right